# Dynamical scaling study for the estimation of dynamical exponent $z$ of three-dimensional XY spin glass model

### v3

@@ Abstract
To analyze the ±$\pm J$ XY spin-glass in three dimensions, we verified a method aimed at obtaining a high-precision dynamical exponent $z$ from the correlation length in the nonequilibrium relaxation process. The obtained $z$ yielded consistent and highly accurate results in previous studies for relatively well-studied models—specifically, the three-dimensional (3D) ferromagnetic Ising model and the 3D ±$\pm J$ Ising model. Building on these previous studies, we used this method and the dynamical scaling method to analyze the 3D ±$\pm J$ XY model and obtained highly precise critical temperatures and exponents. These findings support the spin chirality decoupling picture, explaining the experimental spin-glass phase transition.
I Introduction
@@
The lack of progress in SG research is attributable to the substantial computational time required to achieve equilibrium states as a result of the slow dynamics arising from frustration in SG systems. One effective approach to overcoming this issue is the nonequilibrium relaxation (NER) method. A recent study using this method confirmed the SG universality in the ±$\pm J$ Ising model in three dimensions[2]. In the NER method, obtaining the SG universality expressed by critical exponents $\nu$ necessitates the precise determination of the dynamical critical exponent $z$. The authors of a previous study used the asymptotic behavior of the dynamical Binder parameter $g_{\mathrm{SG}}$ (expressed as $g_{\mathrm{SG}}\sim t^{d/z}$, where $t$ denotes time and $d$ represents the dimensionality of the system) at the critical temperature[3]. By extrapolating the time $t$ to infinity, the authors were able to evaluate $z$ with high precision. However, this method is not applicable when the dynamical Binder parameter $g_{\mathrm{SG}}$ does not exhibit algebraic divergence[4]. One such instance is observed in the Binder parameter for a chiral glass in the 3D ±$\pm J$ XY model. To confirm universality in such SG systems, it remains imperative to compute $z$ with high precision even in such cases.
@@
This paper has two main objectives. The first objective is to investigate the critical phenomena in the ±$\pm J$ XY model, which is an SG system in which the Binder parameter exhibits no algebraic divergence[4]. This approach enables an analysis of critical phenomena in various SG systems, such as Heisenberg systems and gauge glass models. The second objective is to establish the reliability of the estimated $z$ through our study because a precise value of $z$ is needed to study the SG systems through dynamical scaling analysis. To establish the reliability of the estimated $z$, we apply this method to relatively well-studied models—the 3D FM Ising model and the 3D ±$\pm J$ Ising model—and compute $z$.
@@
II Methods
II.1 Models and Observables for Estimating the Critical Temperature
In the present study, we mainly aimed to analyze the ±$\pm J$ XY model in three dimensions. The Hamiltonian can be expressed as
$H=-\sum_{\langle i,j\rangle}J_{ij}\boldsymbol{S}_{i}\cdot\boldsymbol{S}_{j}=-\sum_{\langle i,j\rangle}J_{ij}\cos(\theta_{i}-\theta_{j}).$ (2)
The interaction $J_{ij}$ takes one of two values, $+J$ or $-J$, with the same probability. The summation $\sum_{\langle i,j\rangle}$ is taken over all nearest-neighboring sites. The temperature $T$ is scaled by $J/k_{\mathrm{B}}$. Each spin $\boldsymbol{S}_{i}$ is characterized by an angle $\theta_{i}$ in the $XY$ plane. The linear lattice size is denoted by $L$. The total number of spins is $N=L^{2}(L-1)$, and a skewed periodic boundary condition is imposed.
@@ II.1 Models and Observables for Estimating the Critical Temperature
$\chi_{\rm SG}=\frac{1}{N}\Bigg{[}\sum_{i,j}\langle\boldsymbol{S}_{i}\cdot\boldsymbol{S}_{j}\rangle^{2}\Bigg{]}_{J},$ (3)
[…]
$\langle\boldsymbol{S}_{i}\boldsymbol{S}_{j}\rangle=\frac{1}{n}\sum_{A=1}^{n}\boldsymbol{S}_{i}^{(A)}\boldsymbol{S}_{j}^{(A)},$ (4)
@@
The existence of the chiral-glass (CG) transition in the 3D ±$\pm J$ XY model has been confirmed in numerous previous studies[8, 4, 9]. The CG susceptibility $\chi_{\mathrm{CG}}$ is similarly expressed as
@@
To estimate the critical temperature and some critical exponents of the 3D ±$\pm J$ XY model, we used the following dynamical scaling law for both SG and CG susceptibility[4, 9, 2]:
$\chi(t,T)=t^{\lambda(T)}\Phi\bigg{(}\frac{t}{\tau(T)}\bigg{)},$ (10)
@@
$\lambda(T)=\frac{\gamma}{z_{\mathrm{s}}(T)\nu},$ (11)
$\tau(T)=(T-T_{\rm SG})^{-z_{\mathrm{s}}(T)\nu}.$ (12)
$z_{\mathrm{s}}(T)=z\frac{T_{\mathrm{SG}}}{T}.$ (13)
@@
II.2 Observables and Scaling Law for Estimating the Dynamical Critical Exponent
In the scaling analysis based on the method described in Sec. II.1, the critical temperature $T_{\mathrm{c}}$ and the products of the critical exponents $\gamma/z\nu$ and $z\nu$ can be obtained. Therefore, to estimate all of the critical exponents, it is necessary to calculate $z$. On the basis of Eq. (1), we will explain the method for determining the dynamical correlation length $\xi(t)$ required to calculate $z$.
@@ II.2 Observables and Scaling Law for Estimating the Dynamical Critical Exponent
$f_{\mathrm{SG}}(r)=\Bigg{[}\frac{1}{N}\sum_{i}^{N}\langle\boldsymbol{S}_{i}\cdot\boldsymbol{S}_{i+r}\rangle^{2}\Bigg{]}_{J},$ (14)
@@
$f_{\mathrm{SG}}(r)=\frac{1}{N}\Bigg{[}\sum_{i}^{N}$$\Bigg{\{}\Bigg{(}\frac{1}{n}\sum_{A=1}^{n}\boldsymbol{S}_{i}^{(A)}\cdot\boldsymbol{S}_{i+r}^{(A)}\Bigg{)}^{2}$ (16)
$-\frac{1}{n^{2}}\sum_{A=1}^{n}\Bigg{(}\boldsymbol{S}_{i}^{(A)}\cdot\boldsymbol{S}_{i+r}^{(A)}\Bigg{)}^{2}\Bigg{\}}\Bigg{]}_{J},$
@@
Here, $d$ is the spatial dimension, $\eta_{\mathrm{eff}}$ is an effective scaling exponent, and $f(r,t)$ and $\xi(t)$ are the correlation function and correlation length, respectively, at Monte Carlo step (MCS) $t$. We estimate $\eta_{\mathrm{eff}}$ and dozens of $\xi(t)$ such that all of the $f(r,t)$ data collapse onto a single scaling function $\Phi(\cdot)$. Using the $f(r,t)$ computed at the estimated critical temperature and optimizing the parameters that satisfy the dynamical scaling ansatz Eq. (18), one can obtain sets of $\xi(t)$ at the critical temperature. In conjunction with Eq. (1), this approach enables the value of $z$ to be obtained. When evaluating $z$ in this context, the temperature $T$ in Eq. (13) is set to $T_{\mathrm{SG}}$.
When optimization calculations are performed, dozens of initial quantities $\xi(t)$ are required. For this purpose, we used $\xi(t)$ calculated by the second moment method. The SG and CG correlation length can be calculated by the second moment method as follows:
@@
II.3 Gaussian Process Regression
In this subsection, we briefly explain the method of dynamic scaling analysis based on Gaussian process regression [11, 12, 6, 7]. This method is applied in all scaling analyses, and it is also used in Sec. IV.1 to assess the validity of the scaling plots. Let us write the scaling relation expressed in Eqs. (10) and (18) in the following form:
@@ II.3 Gaussian Process Regression
Here, $E_{i}$ represents the error of $Y_{i}$, and $(\theta_{0},\theta_{1},\theta_{2})$ are the hyperparameters. By maximizing Eq. (22) using an optimization algorithm, the parameters can be estimated. Specifically, the dynamic scaling analysis using Eq. (10) yields the transition temperature and the exponents $\gamma/z\nu$ and $z\nu$, while the analysis using Eq. (18) provides the sets of $\xi(t)$ and $\eta_{\mathrm{eff}}$.
@@
III Confirmation for the Estimation of the Dynamical Exponent
To calculate $z$, we first perform a scaling analysis based on Eq. (18) to obtain a set of dynamical correlation lengths $\xi(t)$ at the transition temperature and then estimate $z$ from the slope of $\xi(t)$ versus $t$ on a double-logarithmic plot according to Eq. (1). However, the reliability of obtained $z$ values using this method has not been thoroughly validated. Therefore, we first apply this method to calculate $z$ for the relatively well-studied 3D FM Ising model and the ±$\pm J$ Ising model to confirm if the results are consistent with previous studies and establish the reliability of this method. Following this confirmation, we apply the method to the primary focus of the present study, the 3D ±$\pm J$ XY model in the next section.
III.1 3D FM Ising Model
To analyze the 3D FM Ising model by the dynamical scaling method, we conducted simulations to 1000 MCS. Within this range, finite-size effects have been confirmed to be absent for $L\geq 159$[5]. We therefore adopted a linear size of $L=301$ in the present study. Under this condition, we computed the correlation function $f(r,t)$ at the critical temperature. Here, the sample number was set to 18,000 and the transition temperature was chosen as the highly accurate value obtained from a previous study, $T_{\mathrm{c}}=4.51152325$[13]. In addition, the MCSs for computing the correlation function were selected to be logarithmically spaced with 10 points between 10 MCS and 1000 MCS.
Figure 1 shows the dynamical correlation function of the 3D FM Ising model at the critical temperature. We subsequently conducted a dynamical scaling analysis based on Eq. (18) to determine the set of correlation lengths required for an accurate estimation of $z$. In an optimization procedure, we discarded short-range correlations ($r<10$) and data of small $f(r,t)$ values ($f(r,t)<10^{-4}$). The initial values of $\{\xi(t)\}$ were set to those of correlations obtained by the second moment method. Figure 2 shows the scaling result based on the dynamical scaling law in Eq. (18). The resultant values of correlation length are plotted in Fig. 3 together with those obtained by the second moment method for comparison. From the inference results and Eq. (1), we evaluated the value of $z$ to be $z=2.018(8)$. In addition, the uncertainty was estimated as the standard deviation of the bootstrap distribution obtained from 10,000 resamplings[14]. This value of $z$ is consistent with the value $z=2.0235(8)$ obtained using the latest reliable methods[15], demonstrating the robustness and accuracy of our scaling analysis based on Eq. (18).
@@ III Confirmation for the Estimation of the Dynamical Exponent
Figure 2: The scaling plot of the dynamical correlation function $f(r,t)$ for the 3D FM Ising model, based on the dynamical scaling law in Eq. (18). The quantities used in the scaling plot are the effective exponent $\eta_{\mathrm{eff}}=0.1433$ and the dozens of correlation length $\{\xi(t)\}$ shown in Fig. 3. These values were obtained through optimization calculations of Gaussian process regression.
@@
Demonstrations of the calculation of $z$ for the 3D ±$\pm J$ Ising model were performed similarly as described in Sec. III.1. Simulations were conducted up to $5\times 10^{5}$ MCS. Within this range, finite-size effects were confirmed to vanish for $L\geq 31$[2]. In the present study, a linear size of $L=61$ was set, $n=256$ replicas were employed, and the sample number was set to 10. The critical temperature was chosen to be the highly accurate value obtained from a previous study, $T_{\mathrm{SG}}=1.178$[2]. The MC steps for computing the correlation function were selected to be logarithmically spaced with 50 points between 100 MCS and $5\times 10^{5}$ MCS.
@@
Figure 4 shows the dynamical SG correlation function of the 3D ±$\pm J$ Ising model at the critical temperature. We subsequently conducted a dynamical scaling analysis based on Eq. (18) to determine the set of correlation lengths required to accurately estimate $z$. During the optimization procedure, short-range correlations ($r<5$) and data with small $f(r,t)$ values ($f(r,t)<5\times 10^{-5}$) were excluded. The initial values of $\{\xi(t)\}$ were set to those of correlations obtained by the second moment method (Eq. (19)). Figure 5 shows the scaling result obtained using the dynamical scaling law in Eq. (18). The resultant values of correlation length are plotted in Fig. 6. From the inference results and Eq. (1), we evaluated the value of $z$ to be $z=5.92(3)$; its uncertainty, defined as the standard deviation of the bootstrap distribution, was assessed using 10 000 bootstrap trials[14]. This result is consistent with the value $z=5.89(3)$ obtained through extrapolation of the dynamical Binder parameter[2], demonstrating the robustness and accuracy of our scaling analysis approach based on Eq. (18).
@@
Figure 4: The dynamical SG correlation function of the 3D ±$\pm J$ Ising model, $f_{\mathrm{SG}}(r,t)$, at the critical temperature $T_{\mathrm{SG}}=1.178$[2] for various Monte Carlo steps $t$.
@@
Figure 5: The scaling plot of the dynamical SG correlation function $f_{\mathrm{SG}}(r,t)$ for the 3D ±$\pm J$ Ising model, based on the dynamical scaling law in Eq. (18). The quantities used in the scaling plot are the effective exponent $\eta_{\mathrm{eff}}=0.1773$ and the dozens of correlation length $\{\xi(t)\}$ shown in Fig. 6. These values were obtained through optimization calculations of Gaussian process regression.
@@
Figure 6: The correlation length $\{\xi(t)\}$, as obtained through optimization calculations of the dynamical scaling law, at the critical point of the ±$\pm J$ Ising model in three dimensions. The dotted line indicates the slope corresponding to the estimated dynamic exponent $z$.
IV Results for 3D ±$\pm J$ XY model

### v4

@@ Abstract
To analyze the $\pm J$ XY spin-glass in three dimensions, we verified a method aimed at obtaining a high-precision dynamical exponent $z$ from the correlation length in the nonequilibrium relaxation process. The obtained $z$ yielded consistent and highly accurate results in previous studies for relatively well-studied models—specifically, the three-dimensional (3D) ferromagnetic Ising model and the 3D $\pm J$ Ising model. Building on these previous studies, we used this method and the dynamical scaling method to analyze the 3D $\pm J$ XY model and obtained highly precise critical temperatures and exponents. These findings support the spin chirality decoupling picture, explaining the experimental spin-glass phase transition.
1 Introduction
@@
The lack of progress in SG research is attributable to the substantial computational time required to achieve equilibrium states as a result of the slow dynamics arising from frustration in SG systems. One effective approach to overcoming this issue is the nonequilibrium relaxation (NER) method. A recent study using this method confirmed the SG universality in the $\pm J$ Ising model in three dimensions[2]. In the NER method, obtaining the SG universality expressed by critical exponents $\nu$ necessitates the precise determination of the dynamical critical exponent $z$. The authors of a previous study used the asymptotic behavior of the dynamical Binder parameter $g_{\mathrm{SG}}$ (expressed as $g_{\mathrm{SG}}\sim t^{d/z}$, where $t$ denotes time and $d$ represents the dimensionality of the system) at the critical temperature[3]. By extrapolating the time $t$ to infinity, the authors were able to evaluate $z$ with high precision. However, this method is not applicable when the dynamical Binder parameter $g_{\mathrm{SG}}$ does not exhibit algebraic divergence[4]. One such instance is observed in the Binder parameter for a chiral glass in the 3D $\pm J$ XY model. To confirm universality in such SG systems, it remains imperative to compute $z$ with high precision even in such cases.
@@
This paper has two main objectives. The first objective is to investigate the critical phenomena in the $\pm J$ XY model, which is an SG system in which the Binder parameter exhibits no algebraic divergence[4]. This approach enables an analysis of critical phenomena in various SG systems, such as Heisenberg systems and gauge glass models. The second objective is to establish the reliability of the estimated $z$ through our study because a precise value of $z$ is needed to study the SG systems through dynamical scaling analysis. To establish the reliability of the estimated $z$, we apply this method to relatively well-studied models—the 3D FM Ising model and the 3D $\pm J$ Ising model—and compute $z$.
@@
2 Methods
2.1 Models and Observables for Estimating the Critical Temperature
In the present study, we mainly aimed to analyze the $\pm J$ XY model in three dimensions. The Hamiltonian can be expressed as
$H=-\sum_{\langle i,j\rangle}J_{ij}\bm{S}_{i}\cdot\bm{S}_{j}=-\sum_{\langle i,j\rangle}J_{ij}\cos(\theta_{i}-\theta_{j}).$ (2)
The interaction $J_{ij}$ takes one of two values, +J+J or J-J, with the same probability. The summation $\sum_{\langle i,j\rangle}$ is taken over all nearest-neighboring sites. The temperature $T$ is scaled by $J/k_{\mathrm{B}}$. Each spin $\bm{S}_{i}$ is characterized by an angle $\theta_{i}$ in the $XY$ plane. The linear lattice size is denoted by $L$. The total number of spins is $N=L^{2}(L-1)$, and a skewed periodic boundary condition is imposed.
@@ 2.1 Models and Observables for Estimating the Critical Temperature
$\chi_{\rm SG}=\frac{1}{N}\Bigg[\sum_{i,j}\langle\bm{S}_{i}\cdot\bm{S}_{j}\rangle^{2}\Bigg]_{J},$ (3)
where $\langle\cdots\rangle$ denotes the thermal average and $\cdots]_{J}$ denotes the average for random-bond configurations. To calculate this quantity, the thermal average is replaced by the average over independent real replicas that consist of different thermal ensembles:
$\langle\bm{S}_{i}\bm{S}_{j}\rangle=\frac{1}{n}\sum_{A=1}^{n}\bm{S}_{i}^{(A)}\bm{S}_{j}^{(A)},$ (4)
@@
The existence of the chiral-glass (CG) transition in the 3D $\pm J$ XY model has been confirmed in numerous previous studies[8, 4, 9]. The CG susceptibility $\chi_{\mathrm{CG}}$ is similarly expressed as
@@
To estimate the critical temperature and some critical exponents of the 3D $\pm J$ XY model, we used the following dynamical scaling law for SG susceptibility[4, 9, 2]:
$\chi_{\mathrm{SG}}(t,T)=t^{\lambda_{\mathrm{SG}}(T)}\Phi\bigg(\frac{t}{\tau_{\mathrm{SG}}(T)}\bigg),$ (10)
@@
$\lambda_{\mathrm{SG}}(T)=\frac{\gamma}{z_{\mathrm{s,SG}}(T)\nu},$ (11)
$\tau_{\mathrm{SG}}(T)=(T-T_{\mathrm{SG}})^{-z_{\mathrm{s,SG}}(T)\nu}.$ (12)
$z_{\mathrm{s,SG}}(T)=z\frac{T_{\mathrm{SG}}}{T}.$ (13)
And for CG case, we used the following dynamical scaling law
$\chi_{\mathrm{CG}}(t,T)=t^{\lambda_{\mathrm{CG}}(T)}\Phi\bigg(\frac{t}{\tau_{\mathrm{CG}}(T)}\bigg),$ (14)
where
$\lambda_{\mathrm{CG}}(T)=\frac{\gamma}{z_{\mathrm{s,CG}}(T)\nu},$ (15)
$\tau_{\mathrm{CG}}(T)=(T-T_{\rm CG})^{-z_{\mathrm{s,CG}}(T)\nu}.$ (16)
$z_{\mathrm{s,CG}}(T)=z\frac{T_{\mathrm{CG}}}{T}.$ (17)
@@
2.2 Observables and Scaling Law for Estimating the Dynamical Critical Exponent
In the scaling analysis based on the method described in Sec. 2.1, the critical temperature $T_{\mathrm{c}}$ and the products of the critical exponents $\gamma/z\nu$ and $z\nu$ can be obtained. Therefore, to estimate all of the critical exponents, it is necessary to calculate $z$. On the basis of Eq. (1), we will explain the method for determining the dynamical correlation length $\xi(t)$ required to calculate $z$.
@@ 2.2 Observables and Scaling Law for Estimating the Dynamical Critical Exponent
$f_{\mathrm{SG}}(r)=\Bigg[\frac{1}{N}\sum_{i}^{N}\langle\bm{S}_{i}\cdot\bm{S}_{i+r}\rangle^{2}\Bigg]_{J},$ (18)
@@
$f_{\mathrm{SG}}(r)=\frac{1}{N}\Bigg[\sum_{i}^{N}$ $\Bigg\{\Bigg(\frac{1}{n}\sum_{A=1}^{n}\bm{S}_{i}^{(A)}\cdot\bm{S}_{i+r}^{(A)}\Bigg)^{2}$ (20)
$-\frac{1}{n^{2}}\sum_{A=1}^{n}\Bigg(\bm{S}_{i}^{(A)}\cdot\bm{S}_{i+r}^{(A)}\Bigg)^{2}\Bigg\}\Bigg]_{J},$
@@
Here, $d$ is the spatial dimension, $\eta_{\mathrm{eff}}$ is an effective scaling exponent, and $f(r,t)$ and $\xi(t)$ are the correlation function and correlation length, respectively, at Monte Carlo step (MCS) $t$. We estimate $\eta_{\mathrm{eff}}$ and dozens of $\xi(t)$ such that all of the $f(r,t)$ data collapse onto a single scaling function $\Phi(\cdot)$.
The concrete procedure for estimating $\xi(t)$ and $z$ in this method is as follows:
  1. 1.
    At the transition temperature, we compute $f(r,t)$ and obtain the initial estimates of $\{\xi(t)\}$ using the second-moment method.
  2. 2.
    Using the obtained $f(r,t)$, we perform an optimization in which the initial values of $\{\xi(t)\}$ are given by the second-moment estimates, and determine the parameters that satisfy the dynamical scaling ansatz Eq. (22).
  3. 3.
    From the optimized $\{\xi(t)\}$, we construct a log–log plot of $\xi(t)$ versus the MCS step $t$, and estimate $z$ from its slope together with Eq. (1).
The second-moment correlation lengths used in the above procedure are given by
@@
2.3 Gaussian Process Regression
In this subsection, we briefly explain the method of dynamic scaling analysis based on Gaussian process regression [11, 12, 6, 7]. This method is applied in all scaling analyses, and it is also used in Sec. 4.1 to assess the validity of the scaling plots. Let us rewrite the scaling relations expressed in Eqs. (10), (14), and (22) in the following form:
@@ 2.3 Gaussian Process Regression
Here, $E_{i}$ represents the error of $Y_{i}$, and $(\theta_{0},\theta_{1},\theta_{2})$ are the hyperparameters. By maximizing Eq. (26) using an optimization algorithm, the parameters can be estimated. Specifically, the dynamic scaling analysis using Eqs. (10) and (14) yield the transition temperature and the exponents $\gamma/z\nu$ and $z\nu$, while the analysis using Eq. (22) provides the sets of $\xi(t)$ and $\eta_{\mathrm{eff}}$.
@@
3 Confirmation for the Estimation of the Dynamical Exponent
To calculate $z$, we first perform a scaling analysis based on Eq. (22) to obtain a set of dynamical correlation lengths $\xi(t)$ at the transition temperature and then estimate $z$ from the slope of $\xi(t)$ versus $t$ on a double-logarithmic plot according to Eq. (1). However, the reliability of obtained $z$ values using this method has not been thoroughly validated. Therefore, we first apply this method to calculate $z$ for the relatively well-studied 3D FM Ising model and the $\pm J$ Ising model to confirm if the results are consistent with previous studies and establish the reliability of this method. Following this confirmation, we apply the method to the primary focus of the present study, the 3D $\pm J$ XY model in the next section.
3.1 3D FM Ising Model
To analyze the 3D FM Ising model by the dynamical scaling method, we conducted simulations up to 5000 MCS. To examine the size dependence in this range, we calculated the relaxation process of the magnetization from the ordered state near the transition temperature. When the size dependence is negligible, the magnetization exhibits a power-law decay behavior[13]. The transition temperature was chosen as the highly accurate value obtained from a previous study, $T_{\mathrm{c}}=4.51152325$[14]. The results are shown in Fig. 1, indicating that the size dependence can be neglected for $L\geq 301$ within this MCS range.
We therefore adopted a linear size of $L=801$ in the present study. Under this condition, we computed the correlation function $f(r,t)$ at the critical temperature. Here, the sample number was set to 1,200. In addition, the MCSs for computing the correlation function were selected to be logarithmically spaced with 10 points between 100 MCS and 5000 MCS.
Figure 1: Size dependence of the magnetization in the 3D ferromagnetic Ising model. The simulations were performed at the transition temperature $T_{\mathrm{c}}=4.51152325$[14]. The number of samples is indicated in the legend. It is found that the size dependence can be neglected for $L\geq 301$.
Figure 2 shows the dynamical correlation function of the 3D FM Ising model at the critical temperature. We subsequently conducted a dynamical scaling analysis based on Eq. (22) to determine the set of correlation lengths required for an accurate estimation of $z$. In an optimization procedure, we discarded short-range correlations ($r<10$) and data of small $f(r,t)$ values ($f(r,t)<10^{-4}$). The initial values of $\{\xi(t)\}$ were set to those of correlations obtained by the second moment method. Figure 3 shows the scaling result based on the dynamical scaling law in Eq. (22). The resultant values of correlation length are plotted in Fig. 4 together with those obtained by the second moment method for comparison. From the inference results and Eq. (1), we evaluated the value of $z$ to be $z=2.038(2)$. In addition, the uncertainty was estimated as the standard deviation of the bootstrap distribution obtained from 100 resamplings[15]. This value of $z$ is compatible with the value $z=2.0346(3)$ obtained using the latest reliable methods[16], demonstrating the robustness and accuracy of our scaling analysis based on Eq. (22).
@@ 3 Confirmation for the Estimation of the Dynamical Exponent
Figure 3: The scaling plot of the dynamical correlation function $f(r,t)$ for the 3D FM Ising model, based on the dynamical scaling law in Eq. (22). The quantities used in the scaling plot are the effective exponent $\eta_{\mathrm{eff}}=0.036(1)$ and the dozens of correlation length $\{\xi(t)\}$ shown in Fig. 4. These values were obtained through optimization calculations of Gaussian process regression.
@@
Demonstrations of the calculation of $z$ for the 3D $\pm J$ Ising model were performed similarly as described in Sec. 3.1. Simulations were conducted up to $5\times 10^{5}$ MCS. Within this range, finite-size effects were confirmed to vanish for $L\geq 31$[2]. In the present study, a linear size of $L=61$ was set, $n=256$ replicas were employed, and the sample number was set to 10. The critical temperature was chosen to be the highly accurate value obtained from a previous study, $T_{\mathrm{SG}}=1.178$[2]. The MC steps for computing the correlation function were selected to be logarithmically spaced with 50 points between 100 MCS and $5\times 10^{5}$ MCS.
@@
Figure 5 shows the dynamical SG correlation function of the 3D $\pm J$ Ising model at the critical temperature. We subsequently conducted a dynamical scaling analysis based on Eq. (22) to determine the set of correlation lengths required to accurately estimate $z$. During the optimization procedure, short-range correlations ($r<5$) and data with small $f(r,t)$ values ($f(r,t)<5\times 10^{-5}$) were excluded. The initial values of $\{\xi(t)\}$ were set to those of correlations obtained by the second moment method (Eq. (23)). Figure 6 shows the scaling result obtained using the dynamical scaling law in Eq. (22). The resultant values of correlation length are plotted in Fig. 7. From the inference results and Eq. (1), we evaluated the value of $z$ to be $z=5.895(7)$; its uncertainty, defined as the standard deviation of the bootstrap distribution, was assessed using 100 bootstrap trials[15]. This result is consistent with the value $z=5.89(3)$ obtained through extrapolation of the dynamical Binder parameter[2], demonstrating the robustness and accuracy of our scaling analysis approach based on Eq. (22).
@@
Figure 5: The dynamical SG correlation function of the 3D $\pm J$ Ising model, $f_{\mathrm{SG}}(r,t)$, at the critical temperature $T_{\mathrm{SG}}=1.178$[2] for various Monte Carlo steps $t$.
@@
Figure 6: The scaling plot of the dynamical SG correlation function $f_{\mathrm{SG}}(r,t)$ for the 3D $\pm J$ Ising model, based on the dynamical scaling law in Eq. (22). The quantities used in the scaling plot are the effective exponent $\eta_{\mathrm{eff}}=-0.183(2)$ and the dozens of correlation length $\{\xi(t)\}$ shown in Fig. 7. These values were obtained through optimization calculations of Gaussian process regression.
@@
Figure 7: The correlation length $\{\xi(t)\}$, as obtained through optimization calculations of the dynamical scaling law, at the critical point of the $\pm J$ Ising model in three dimensions. The dotted line indicates the slope corresponding to the estimated dynamic exponent $z$.
4 Results for 3D $\pm J$ XY model